OpenStudy (anonymous):

need help! use substitution method 2x+6y=4 -3x+y=24...answer must be an ordered pair...Thanks!

jhonyy9 (jhonyy9):

y=24+3x

OpenStudy (anonymous):

x=2-3y -3(2-3y)=24 -6+9y=24 9y=20 y=20/9 x=2-3(20/9} x=-14/3

OpenStudy (anonymous):

nevermind i missed a y

OpenStudy (anonymous):

y=24+3x 2x+6(24+3x)=4 2x+144+18x=4 20x=-140 x=-7 y=24+3(-7) y=3

OpenStudy (amistre64):

2 6 = 4 -3 1 = 24 1 -3 = 24 0 10 = -70 1 -3 = 24 0 1 = -7 1 0 = 3 0 1 = -7 y x maybe

OpenStudy (amistre64):

i spose i coulda forgone a step and just reduced the top first :)

OpenStudy (ybarrap):

2x + 6y = 4 (*) -3x + y = 24 (**) y = 24 + 3x, from (**) above = 3*(8 + x) (***) 2x + 6*3*(8 + x) = 4, substituting (***) into (*) above 2x + 18*8 + 18x = 4 20x + 144 = 4 20x = 4 - 144 = -140 20x/20 = -140/20 x = -7*20/20 = -7 (****) y = 24 + 3*(-7) substituting (****) into (**) above = 24 - 21 = 3 Ans. (-7,3)